OpenStudy (anonymous):

Two cars drive in the same direction along a straight road, at a constant speed - one at 70 km/hr and the other at 80 km/hr. Assuming that they start at the same point and at the same time, how much sooner would the faster car arrive at a destination 460 km away? a. 49 minutes c. 5.75 hrs b. 6.6 hrs d. 741 minutes

OpenStudy (anonymous):

d. 741 minutes

OpenStudy (anonymous):

only if your assumption of them starting at the same point and time is correct if its wrong then we dont know....

OpenStudy (anonymous):

Joe is on a cross-country trip from New York to Los Angeles. He drives for 45.0 min at 70.0 km/h, 10.0 min at 95 km/h, and 30.0 min at 55.0 km/h, and he spends 20.0 min eating lunch and buying gas. What is the average speed for this part of his trip? a. 67.6 km/h c. 54.7 km/h b. 68.4 km/h d. 53.8 km/h @soty2013

OpenStudy (anonymous):

D.) 741 minutes is not correct.

OpenStudy (anonymous):

then what is it?

OpenStudy (anonymous):

460 / 80 = 5.75 hours 460 / 70 = 6.5714285714285714285714285714286 hours 6.5714285714285714285714285714286 - 5.75 = 0.82142857142857142857142857142857 0.82142857142857142857142857142857 * 60 = 49.285714285714285714285714285714 minutes. The answer is A.